OpenStudy (anonymous):

find the equation of a line that satisfies the given conditions; x-intercept 2, and y-intercept -8

OpenStudy (anonymous):

so the points are (2,0) and (0,-8) use the definition of slope which is \[m = \frac{y2 - y1}{x2 - x1}\] to find the slope and then use the point slope formula \[y - y1 = m(x - x1)\]

OpenStudy (anonymous):

To write the equation

OpenStudy (anonymous):

So what do you get for the slope?

OpenStudy (anonymous):

slope is 4

OpenStudy (anonymous):

yeah the slope is 4

OpenStudy (anonymous):

Good.. so m = 4 and we can choose anyone of those points above to write the equation.. So pick what point you want to use.

OpenStudy (anonymous):

y+8=4x